Write the standard form of the equation of the line with a slope 2/7 passing through the point (7, -4)
A line (L1) has a slope of 1. The line through which of the following points is perpendicular to L1?

PLEASE HELP ME!

Answer by nerdybill(7384)

Write the standard form of the equation of the line with a slope 2/7 passing through the point (7, -4)
.
plug into "point-slope" form:
y - y1 = m(x - x1)
y - (-4) = (2/7)(x - 7)
y + 4 = (2/7)x - 2
y = (2/7)x - 6
.
A line (L1) has a slope of 1. The line through which of the following points is perpendicular to L1?
you didn't provide the points...